OpenStudy (anonymous):

Medal Solve sqrtx-1=x-3 check for extraneous solutions A. No solution B. x=5 C. x=2 D. x=2, 5

OpenStudy (neer2890):

is it \[\sqrt{x}-1=x-3\] or \[\sqrt{x-1}=x-3\]??

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Square both sides. then you will get a quadratic equation. Solution to Q.E of the form is \[a^{2}+bx +c=0 \] \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (neer2890):

joseph91 is right

OpenStudy (neer2890):

x-1=(x-3)^2 x-1=x^2-6x+9 \[x ^{2}-7x+10=0\]

OpenStudy (neer2890):

solve this equation there are two roots of this equation (2,5) (x-2)(x-5)=0

OpenStudy (neer2890):

quadratic equation should be \[ax ^{2}+bx+c=0\] joseph by mistaken wrote \[a ^{2}\] instead of\[ax ^{2}\]

OpenStudy (anonymous):

sorry for mistake